# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 6 Learning in Recurrent Models
Figure 17: Overview of the organization of Section 6.
Backpropagation Through Time (BPTT, see Section 2) is the de-facto standard algorithm for training recurrent models. It involves unrolling (i.e., virtually replicating) a recurrent network over the whole input sequence of length $L$, sharing the same parameters $L$ times, and “backpropagating” the error from the $L$-th instance to the first one. In Section 2 we emphasized the advantages and drawbacks of BPTT, such as ($i$) learning issues due to vanishing/exploding gradients and ($ii$) the high memory and computational requirements, that hinder the ability to handle long range sequences. Indeed, whenever data come in the form of a stream [93], the BPTT requirements make it a non-feasible choice for learning online on a potentially infinite sequence, given the difficulties in unrolling the network over long time horizons. Most of the models described in previous sections propose to alleviate such issues by careful architectural designs, but still keeping BPTT as the learning procedure, and assuming that the whole sequence is available beforehand. In this section, we overview recent alternative learning mechanisms that try to tackle such drawbacks ($i$ and $ii$) of BPTT. We leverage a more general form of the recurrent mechanism described in Eq. (2), explicitly including a state transition function $F(\cdot)$ and an output readout function $G(\cdot)$, defined as follows (to simplify the notation when describing the approaches we review),
@@
$=J_{t}M_{t-1}+\bar{M}_{t},$
where $J_{t}=\frac{\partial x_{t}}{\partial x_{t-1}}$ is the Jacobian of the actual state w.r.t. the previous state and $\bar{M}_{t}=\frac{\partial x_{t}}{\partial\theta^{F}_{t}}$ is called the immediate influence. Notice the distinction between $\theta^{F}_{s}$ and $\theta^{F}_{t}$: the former is about instances of the weights in the previous time instants, while the latter is about the current weight values. It is possible to obtain the derivatives of the loss function with respect to $\theta^{F}$ by,
@@
where $\bar{c}_{t}=\frac{\partial\ell_{t}}{\partial x_{t}}$ is called the immediate credit assignment vector. Differently form the just described case of $\theta^{F}$, it is natural to directly learn $\theta^{G}$ online, because only information at present time $t$ is required to calculate the gradient $\frac{\partial\ell_{t}}{\partial\theta^{G}}$. RTRL suggests to propagate the partials $\frac{\partial x_{t}}{\partial x_{t-1}}$ and $\frac{\partial x_{t}}{\partial\theta^{F}}$ from timestep $t$ to $t+1$. This is based on the intuition that there is significant overlap in the product term (see Eq. (5) and Eq. (6)) from time $t$ to $t+1$, allowing for recursive computation. This algorithm is online and past-facing, which is defined by the fact that only previously computed quantities are used in the computations [184]. These properties can be directly observed in the above formulas, which explicitly depend only on timesteps $t$ and $t-1$. RTRL is also deterministic and provides the solution in closed form. Since RTRL requires, for each time step, the storage of $M_{t}$, which involves the gradients of each component of the state (${d_{\text{s}}}$ elements) with respect to all the parameters involved in the state computation (i.e., $|\theta^{F}|={d_{\text{s}}}\cdot{d_{\text{in}}}+{d_{\text{s}}}^{2}$), its memory complexity is $\mathcal{O}({d_{\text{s}}}({d_{\text{s}}}\cdot{d_{\text{in}}}+{d_{\text{s}}}^{2}))\sim\mathcal{O}({d_{\text{s}}}^{3})$ and, for the computation of $J_{t}M_{t-1}$, its time complexity is $\mathcal{O}({d_{\text{s}}}^{4})$. Early attempts suffered large memory overhead limiting its usage, and while recent attempts [279, 184] have been more successful, these methods still fall short of BPTT performance, and so trainability of RNNs is still a significant issue. Several other early attempts to solve the shortcomings of BPTT were motivated by the human way to learn from perceptual stimuli, which are intrinsically continuous over time and not pre-buffered finite-length sequences randomly shuffled to cope with stochastic gradient descent [288]. Thus, from [302], the proposal of “an on-line algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream is required”. Even the LSTMs [105] were introduced with a learning algorithm that unlinke full BPTT is “local in space and time”, where “there is no need to store activation values observed during sequence processing in a stack with potentially unlimited size”. Recurrent Backpropagation (RBP) [1, 226, 221] avoids the need to unroll the entire forward pass, as required by BPTT, by directly computing the gradient w.r.t. the learnable parameters at a steady state of Eq. (63), exploiting the implicit function theorem [246] and achieving constant memory complexity w.r.t. the number of processing steps. In details, RBP assumes that the dynamics of the state transition $F(\cdot)$ reach an equilibrium with a steady-state hidden state $x^{*}$, i.e., $x^{*}=F(x^{*},u,\theta^{F})$ when processing an input $u$, which is fixed and not time-dependent. [^20] RBP formulation assumes a fixed not time-dependent input $u$. However, in common scenarios where data is i.i.d, the sequential input data can be interpreted as sampled from a stationary distribution. As a consequences, RBP can be applied, since the steady state holds in expectation[162]. In this condition, it is possible to construct a function $\Psi(x,\theta^{F})=x-F(x,u,\theta^{F})$ such that, when the system dynamic has reached the equilibrium (i.e., a fixed point of the state, $x^{*}$), $\Psi(x,\theta^{F})=0$. Differentiating $\Psi(x,\theta^{F})$ w.r.t the parameters $\theta^{F}$ at $x^{*}$ and rearranging the terms yields the gradient of the steady state $x^{*}$ w.r.t. the parameters of the stable dynamical system,
@@
via fixed point iterations. RBP was the learning algorithm exploited in the seminal works introducing Graph Neural Networks (GNNs) [80, 256], which can be considered as generalizations of RNNs to handle graph-structured input data [274, 70, 193]. Indeed, the current literature refers to the GNNs by Scarselli et al. [256] as Recurrent GNNs (RecGNNs) [307]. Inference in RecGNNs can be interpreted as a diffusion process along the graph up to the convergence onto fixed points of the nodal states. RecGNNs ensure the RBP conditions ($i$) and ($ii$) by forcing $F(\cdot)$ to be a contraction map on Banach space, a choice that however poses some strong limitations on the model capacity and that is difficult to satisfy for general neural networks. When used with models that satisfy such assumptions, the main computational cost of RBP lies in solving a linear system (i.e., where the most expensive operation is the matrix-vector product $J_{F,x^{*}}^{T}z$) which has constant memory and computation time w.r.t. the number of unrolling steps.
@@
If we indicate with $s$ the level of sparsity of the matrix $M_{t}$, and we define $d$ as $d:=1-s$, the SnAp-1 algorithm ($n=1$) for a fully connected RNN has a memory complexity of $\mathcal{O}({d_{\text{s}}}+d|\theta^{F}|)$ and a time complexity of $O(d({d_{\text{s}}}^{2}+|\theta^{F}|))$. RTRL can be also made computationally tractable introducing different RNN structures and adopting specific learning processes exploiting these architectures for enabling scalable, unbiased and noise-free gradient estimation. For example, Schmidhuber et al. [117] proposed to apply RTRL gradient computation to RNNs with element-wise recurrence (eLSTM). Under this architectural assumption, it is possible to obtain forward recursive formulas for the inference matrix, which have a memory complexity of $\mathcal{O}({d_{\text{in}}}{d_{\text{s}}})$ and a per-step time complexity of $\mathcal{O}({d_{\text{s}}}^{2})$. Always focusing on the net structure, Sutton et al. proposed Columnar Neural networks (Col-NNs) [123], Constructive Networks and Columnar-Constructive networks (CCNs) [121, 122]. In Col-NNs [123], the network structure is restricted to be composed of independent, potentially deep columns. Each column presents a scalar recurrent state, which is not connected to the states of the other columns. Therefore, it is possible to apply RTRL to each column individually reducing the computational cost to $\mathcal{O}(|\theta_{F}|)$, which means that RTRL for Col-NNs scales linearly in the size of the parameters. However, the structure of Col-NNs lacks hierarchical recurrent features. To introduce this hierarchy, Constructive Networks [121] have been introduced, learning the recurrent network one feature at a time. The learning process prioritizes the acquisition of weights associated with the first recurrent feature, exclusively linked to the input. Subsequently, these weights are frozen, facilitating progression to the subsequent hidden unit. This unit can now establish connections with preceding recurrent features. Notably, the output weights remain dynamic, undergoing continual updates. By sequentially focusing on discrete subsets of the network during training, Constructive Networks incur even lower per-step computations than Columnar Networks. Consequently, RTRL can be implemented efficiently. Constructive Networks have the limitation to be incapable to learn multiple features in parallel. To overcome this issue, CCNs [122] learn multiple columns that take as input the features of all the existing frozen columns, iteratively constructing multiple columns of recurrent features in parallel. For other approximations of RTRL, see [184]. Finally inspired by the success of networks with recurrent units [213] and leveraging on the property that recurrence with only self-loops greatly simplifies RTRL [200, 198] in [320] the authors propose a modification of RTRL that is tailored on architectures with linear recurrent blocks interconnected through nonlinear layers. In particular, within this setting, they propose an approximation of the gradient in case where more than one recurrent layer is stacked, artificially reducing the dependencies over time of the multiple layers.
@@
[…]
where $\mathcal{L}$ is the loss function of Eq. (4). Thus, such models are referred to as “lifted” because the usual parameter search space of $\theta$-variables is lifted to an higher dimensional space composed by $(\theta,x)$-variables. The common approach is to transform the non-smooth constrained optimization problem into a smooth unconstrained problem in the enlarged space. Early approaches were proposed for feed-forward architectures (but can be easily extended to RNNs) [32], by adding quadratic penalties to approximately enforce the equality constraints. Succeeding works [284] use Lagrange multipliers to exactly enforce equality constraints and optimize via Alternating Direction Method of Multipliers (ADMM) and Bregman iteration. A clear advantage of these methods is the ability to decompose the training problem into multiple, local sub-problems which can be solved efficiently. Recently, works by Askari et al. [7, 92] proposed convex/biconvex formulations that can be optimized using Block Coordinate Descent, that were extended to RNNs in [254]. Marra et al. [183] investigated the connections of lifted methods to Backpropagation, and proposed an hard-constraining scheme, referred to as Local Propagation, based on the augmented Lagrangian. Learning consists in a differential optimization problem converging towards a saddle point of the Lagrangian. Interestingly, this approach has been extended to devise a novel learning algorithm for RecGNNs [286] and allowing to exploit deep RecGNNs [285, 175], by implicitly expressing the state convergence procedure via a constraint satisfaction mechanism. This removes the need of iterative procedures to be applied at each training epochs, and the network unfolding of the original model, avoiding the harsh constraints imposed by RBP. Despite achieving good performances in classic benchmarks, the memory requirements of lifted methods, due to the introduction of additional trainable parameters, is their main drawback.
@@
¯$\overline{\ell}(\theta)$ $:=\ell_{t}(\theta)+\frac{\alpha}{2}\|\theta-\overline{\theta}_{t}-\frac{1}{2\alpha}\nabla\overline{\ell}_{t-1}(\theta_{t})\|^{2},$ (75)
$\hskip-2.84544pt\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\overline{\ell}(\theta)\big{|}_{\theta=\theta_{t}},$
$\hskip-2.84544pt\overline{\theta}_{t+1}=\frac{1}{2}(\overline{\theta}_{t}+\theta_{t+1})-\frac{1}{2\alpha}\nabla\ell_{t}(\theta_{t+1}),$
@@
where $\alpha$ is a weighing factor, ¯$\overline{\ell}(\theta)$ denotes the augmented instantaneous risk function and $\overline{\theta}$ is the state-vector that summarizes past losses. Indeed, such a “summary” represents a running average of past $\theta_{t}$ plus a correction term, that enforces stability in updates and convergence of the parameters toward a fixed point (i.e., only in this case hidden state trajectories simulate the usual static time invariant RNN parameters). FPTT requires $\mathcal{O}(L)$ gradient computations for an $L$-length sequence, while BPTT computes gradient only once, when the whole sequence has been processed. However, the constants (when evaluating the complexity) involved in taking gradient for the full-sequence are higher than computing single-step gradients. From the memory point of view, BPTT stores intermediate hidden state about the whole sequence, i.e., $\mathcal{O}(L)$, while FPTT does not require storing hidden states, i.e., it is $\mathcal{O}(1)$. The authors of FPTT additionally propose a more efficient variant that performs updates restricted to $K$ time-steps, referred to as FPTT-K. Given that Eq. (75) exploits instantaneous loss computations, when the task does not involve step-wise supervisions (e.g., terminal prediction), FPTT leverages an alternative formulation that approximate the previous one [130]. Going beyond FPTT, another alternative method for addressing the exploding-vanishing gradient problem in processing long-sequences, is the one of Park et al. [217]. Such a method consists in a novel initialization scheme for RNNs enhancing learning of temporal dynamics. In order to achieve a stable limit cycle, defined as an attracting ring manifold where the neural activations form a periodic trajectory, the authors structure the weight matrix using a scaled rotation matrix. This results in a block orthogonal matrix arrangement where, within each $2\times 2$ block, the behavior of the paired neurons exhibits spontaneous oscillations. The exploding-vanishing gradient issue can also be related to the presence of bifurcations in the parameter space of the RNN, which represent qualitative shifts in network dynamics due to parameter variations. As shown by Eisenmann et al. [63], bifurcations are associated with abrupt changes in stability regions and the topological structure of the state space. The authors of [63] propose a heuristic algorithm (SCYFI, which stands for Searcher for Cycles and Fixed points) for identifying fixed points and cycles in ReLU-based RNNs and determining their existence and stability regions in the parameter space, along with eventual bifurcations. Once that fixed points, cycles and bifurcations have been identified, Generalized Teacher Forcing (GTF) (a method aimed at redirecting diverging trajectories towards their intended targets) tends to circumvent bifurcations during training. Finally, we mention the different route followed by Echo State Networks and, more generally, by instances of Reservoir Computing, where a large (usually non-learned) sparse connectivity layer is followed by a trainable readout function. The reader can find more information in [255, 194, 71].

### v2

@@ 6 Learning in Recurrent Models
Figure 17: Conceptual overview of the organization of Section 6, which classifies learning approaches for recurrent models based on methodological principles. Yellow boxes denote thematic categories, and representative methods are listed in gray.
Backpropagation Through Time (BPTT, see Section 2) is the de-facto standard algorithm for training recurrent models. It involves unrolling (i.e., virtually replicating) a recurrent network over the whole input sequence of length $L$, sharing the same parameters $L$ times, and “backpropagating” the error from the $L$-th instance to the first one. In Section 2, we have emphasized the advantages and drawbacks of BPTT, such as ($i$) learning issues due to vanishing/exploding gradients and ($ii$) the high memory and computational requirements. These requirements hinder the ability to handle long-range sequences. Indeed, whenever data come in the form of a stream [193], the BPTT requirements make it a non-feasible choice for online learning on a potentially infinite sequence, given the difficulties in unrolling the network over long time horizons. Most of the models described in previous sections propose to alleviate such issues by careful architectural designs, but still retain BPTT as the learning procedure, and assume that the whole sequence is available beforehand.
In this section, we overview recent alternative learning mechanisms that aim to address such drawbacks ($i$ and $ii$) of BPTT.
We leverage a more general form of the recurrent mechanism described in Eq. (2), explicitly including a state transition function $F(\cdot)$ and an output readout function $G(\cdot)$, defined as follows. This general form simplifies the notation in the subsequent descriptions of the reviewed approaches.
@@
$=J_{t}M_{t-1}+\bar{M}_{t}.$
We now recall that the Jacobian matrix of a vector-valued function is a matrix of all the first-order partial derivatives of the function, w.r.t. its arguments. More precisely, let
$f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m},\quad f(x)=\begin{bmatrix}f_{1}(x_{1},x_{2},\dots,x_{n})\\ f_{2}(x_{1},x_{2},\dots,x_{n})\\ \vdots\\ f_{m}(x_{1},x_{2},\dots,x_{n})\end{bmatrix}$. (66)
Then, the Jacobian matrix of $f$ with respect to $x$ is defined as:
$J_{f}(x)=\begin{bmatrix}\frac{\partial f_{1}}{\partial x_{1}}&\frac{\partial f_{1}}{\partial x_{2}}&\cdots&\frac{\partial f_{1}}{\partial x_{n}}\\ \frac{\partial f_{2}}{\partial x_{1}}&\frac{\partial f_{2}}{\partial x_{2}}&\cdots&\frac{\partial f_{2}}{\partial x_{n}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial f_{m}}{\partial x_{1}}&\frac{\partial f_{m}}{\partial x_{2}}&\cdots&\frac{\partial f_{m}}{\partial x_{n}}\end{bmatrix}$. (67)
In Eq. (65), $J_{t}=\frac{\partial x_{t}}{\partial x_{t-1}}$ is the Jacobian of the actual state w.r.t. the previous state and $\bar{M}_{t}=\frac{\partial x_{t}}{\partial\theta^{F}_{t}}$ is called the immediate influence. Notice the distinction between $\theta^{F}_{s}$ and $\theta^{F}_{t}$: the former is about instances of the weights in the previous time instants, while the latter is about the current weight values. It is possible to obtain the derivatives of the loss function with respect to $\theta^{F}$ by,
@@
where $\bar{c}_{t}=\frac{\partial\ell_{t}}{\partial x_{t}}$ is called the immediate credit assignment vector. Differently form the just described case of $\theta^{F}$, it is natural to directly learn $\theta^{G}$ online, because only information at present time $t$ is required to calculate the gradient $\frac{\partial\ell_{t}}{\partial\theta^{G}}$. RTRL suggests to propagate the partials $\frac{\partial x_{t}}{\partial x_{t-1}}$ and $\frac{\partial x_{t}}{\partial\theta^{F}}$ from timestep $t$ to $t+1$. This is based on the intuition that there is significant overlap in the product term (see Eq. (5) and Eq. (6)) from time $t$ to $t+1$, allowing for recursive computation. This algorithm is online and past-facing, which is defined by the fact that only previously computed quantities are used in the computations [45]. These properties can be directly observed in the above formulas, which explicitly depend only on timesteps $t$ and t1t-1. RTRL is also deterministic and provides the solution in closed form. Since RTRL requires, for each time step, the storage of $M_{t}$, which involves the gradients of each component of the state ($d_{\text{s}}$} elements) with respect to all the parameters involved in the state computation (i.e., $|\theta^{F}|={d_{\text{s}}}\cdot{d_{\text{in}}}+{d_{\text{s}}}^{2}$), its memory complexity is $\mathcal{O}({d_{\text{s}}}({d_{\text{s}}}\cdot{d_{\text{in}}}+{d_{\text{s}}}^{2}))\sim\mathcal{O}({d_{\text{s}}}^{3})$ and, for the computation of $J_{t}M_{t-1}$, its time complexity is $\mathcal{O}({d_{\text{s}}}^{4})$. Early attempts suffered large memory overhead limiting its usage, and while recent attempts [194, 45] have been more successful, these methods still fall short of BPTT performance, and so trainability of RNNs is still a significant issue. Several other early attempts to solve the shortcomings of BPTT were motivated by the human way to learn from perceptual stimuli, which are intrinsically continuous over time and not pre-buffered finite-length sequences randomly shuffled to cope with stochastic gradient descent [195]. Thus, from [20], the proposal of “an on-line algorithm, designed to be used to train a network while it runs; no manual state resets or segmentations of the training stream is required”. Even the LSTMs [28] were introduced with a learning algorithm that unlinke full BPTT is “local in space and time”, where “there is no need to store activation values observed during sequence processing in a stack with potentially unlimited size”. Recurrent Backpropagation (RBP) [196, 197, 198] avoids the need to unroll the entire forward pass, as required by BPTT, by directly computing the gradient w.r.t. the learnable parameters at a steady state of Eq. (63), exploiting the implicit function theorem [199] and achieving constant memory complexity w.r.t. the number of processing steps. In details, RBP assumes that the dynamics of the state transition $F(\cdot)$ reach an equilibrium with a steady-state hidden state $x^{*}$, i.e., $x^{*}=F(x^{*},u,\theta^{F})$ when processing an input $u$, which is fixed and not time-dependent. [^21] RBP formulation assumes a fixed not time-dependent input $u$. However, in common scenarios where data is i.i.d, the sequential input data can be interpreted as sampled from a stationary distribution. As a consequences, RBP can be applied, since the steady state holds in expectation[200]. In this condition, it is possible to construct a function $\Psi(x,\theta^{F})=x-F(x,u,\theta^{F})$ such that, when the system dynamic has reached the equilibrium (i.e., a fixed point of the state, $x^{*}$), $\Psi(x,\theta^{F})=0$. Differentiating $\Psi(x,\theta^{F})$ w.r.t the parameters $\theta^{F}$ at $x^{*}$ and rearranging the terms yields the gradient of the steady state $x^{*}$ w.r.t. the parameters of the stable dynamical system,
@@
via fixed point iterations. RBP was the learning algorithm exploited in the seminal works introducing Graph Neural Networks (GNNs) [201, 202], which can be considered generalizations of RNNs to handle graph-structured input data [203, 204, 205]. Indeed, the current literature refers to GNNs by Scarselli et al. [202] as Recurrent GNNs (RecGNNs) [11]. Inference in RecGNNs can be interpreted as a diffusion process along the graph up to the convergence to fixed points of the nodal states. RecGNNs ensure the RBP conditions ($i$) and ($ii$) by forcing $F(\cdot)$ to be a contraction map on a Banach space, a choice thatneverthelessposes some strong limitations on the model capacity , and is difficult to satisfy. When used with models that satisfy such assumptions, the main computational cost of RBP lies in solving a linear system (i.e., where the most expensive operation is the matrix-vector product $J_{F,x^{*}}^{T}z$), which has constant memory and computation time with respect to the number of unrolling steps.
@@
If we indicate with $s$ the level of sparsity of the matrix $M_{t}$, and we define $d$ as d:=1sd:=1-s, the SnAp-1 algorithm ($n=1$) for a fully connected RNN has a memory complexity of $\mathcal{O}({d_{\text{s}}}+d|\theta^{F}|)$ and a time complexity of $O(d({d_{\text{s}}}^{2}+|\theta^{F}|))$. RTRL can be also made computationally tractable introducing different RNN structures and adopting specific learning processes exploiting these architectures for enabling scalable, unbiased and noise-free gradient estimation. For example, Schmidhuber et al. [207] proposed to apply RTRL gradient computation to RNNs with element-wise recurrence (eLSTM). Under this architectural assumption, it is possible to obtain forward recursive formulas for the inference matrix, which have a memory complexity of $\mathcal{O}({d_{\text{in}}}{d_{\text{s}}})$ and a per-step time complexity of $\mathcal{O}({d_{\text{s}}}^{2})$. Always focusing on the net structure, Sutton et al. proposed Columnar Neural networks (Col-NNs) [208], Constructive Networks and Columnar-Constructive networks (CCNs) [209, 27]. In Col-NNs [208], The network structure is restricted to be composed of independent, potentially deep columns. Each column presents a scalar recurrent state, which is not connected to the states of the other columns. Therefore, it is possible to apply RTRL to each column individually, reducing the computational cost to $\mathcal{O}(|\theta_{F}|)$, which means that RTRL for Col-NNs scales linearly with the size of the parameters. However, the structure of Col-NNs lacks hierarchical recurrent features. To introduce this hierarchy, Constructive Networks [209] have been introduced, learning the recurrent network one feature at a time. The learning process prioritizes the acquisition of weights associated with the first recurrent feature, exclusively linked to the input. Subsequently, these weights are frozen, facilitating progression to the subsequent hidden unit. This unit can now establish connections with preceding recurrent features. Notably, the output weights remain dynamic, undergoing continual updates. By sequentially focusing on discrete subsets of the network during training, Constructive Networks incur even lower per-step computations than Columnar Networks. Consequently, RTRL can be implemented efficiently. Constructive Networks have the limitation of being unable to learn multiple features in parallel. To overcome this issue, CCNs [27] learn multiple columns that take as input the features of all the existing frozen columns, iteratively constructing multiple columns of recurrent features in parallel. For other approximations of RTRL, see [45]. Finally, inspired by the success of networks with recurrent units [43] and leveraging the fact that recurrence with only self-loops greatly simplifies RTRL [210, 211], in [212] the authors propose a modification of RTRL that is tailored to architectures with linear recurrent blocks interconnected through nonlinear layers. In particular, within this setting, they propose an approximation of the gradient in the case where more than one recurrent layer is stacked, artificially reducing the dependencies over time across layers.
@@
$\text{s.t.}\qquad x_{t}=F(x_{t-1},u_{t},\theta^{F}),$
where $\mathcal{L}$ is the loss function of Eq. (4). Thus, such models are referred to as “lifted” because the usual parameter search space of $\theta$-variables is lifted to an higher dimensional space composed by $(\theta,x)$-variables. The common approach is to transform the non-smooth constrained optimization problem into a smooth unconstrained problem in the enlarged space. Early approaches were proposed for feed-forward architectures (but can be easily extended to RNNs) [215], by adding quadratic penalties to approximately enforce the equality constraints. Succeeding works [216] use Lagrange multipliers to exactly enforce equality constraints and optimize via the Alternating Direction Method of Multipliers (ADMM) and Bregman iteration. A clear advantage of these methods is the ability to decompose the training problem into multiple, local sub-problems, which can be solved efficiently. Recently, works by Askari et al. [217, 218] proposed convex and biconvex formulations that can be optimized using Block Coordinate Descent. These methods were extended to RNNs in [217]. Marra et al. [219] investigated the connections of lifted methods to Backpropagation, and proposed a hard-constraining scheme, referred to as Local Propagation, based on the augmented Lagrangian. Learning consists of a differential optimization problem converging towards a saddle point of the Lagrangian. Interestingly, this approach has been extended to devise a novel learning algorithm for RecGNNs [220], allowing for the exploitation of deep RecGNNs [221, 222], by implicitly expressing the state convergence procedure via a constraint satisfaction mechanism. This removes the need for iterative procedures to be applied at each training epoch, and eliminates the network unfolding of the original model and the harsh constraints imposed by RBP. Despite achieving good performance in classic benchmarks, the memory requirements of lifted methods, due to the introduction of additional trainable parameters, remain their main drawback.
@@
$\overline{\ell}(\theta)$ $=\ell_{t}(\theta)+\frac{\alpha}{2}\|\theta-\overline{\theta}_{t}-\frac{1}{2\alpha}\nabla\overline{\ell}_{t-1}(\theta_{t})\|^{2},$ (77)
$\thinspace\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\overline{\ell}(\theta)\big{|}_{\theta=\theta_{t}},$
$\thinspace\overline{\theta}_{t+1}=\frac{1}{2}(\overline{\theta}_{t}+\theta_{t+1})-\frac{1}{2\alpha}\nabla\ell_{t}(\theta_{t+1}),$
@@
where $\alpha$ is a weighing factor, $\overline{\ell}(\theta)$ denotes the augmented instantaneous risk function and $\overline{\theta}$ is the state-vector that summarizes past losses. Indeed, such a “summary” represents a running average of past $\theta_{t}$ plus a correction term, that enforces stability in updates and convergence of the parameters toward a fixed point (i.e., only in this case hidden state trajectories simulate the usual static time invariant RNN parameters). FPTT requires $\mathcal{O}(L)$ gradient computations for an $L$-length sequence, while BPTT computes gradient only once, when the whole sequence has been processed. However, the constants (when evaluating the complexity) involved in taking gradient for the full-sequence are higher than computing single-step gradients. From the memory point of view, BPTT stores intermediate hidden state about the whole sequence, i.e., $\mathcal{O}(L)$, while FPTT does not require storing hidden states, i.e., it is $\mathcal{O}(1)$. The authors of FPTT additionally propose a more efficient variant that performs updates restricted to $K$ time-steps, referred to as FPTT-K. Given that Eq. (77) exploits instantaneous loss computations, when the task does not involve step-wise supervisions (e.g., terminal prediction), FPTT leverages an alternative formulation that approximate the previous one [58].
Going beyond FPTT, another alternative method for addressing the exploding-vanishing gradient problem in processing long-sequences, is the one of Park et al. [223]. Such a method consists in a novel initialization scheme for RNNs enhancing learning of temporal dynamics. In order to achieve a stable limit cycle, defined as an attracting ring manifold where the neural activations form a periodic trajectory, the authors structure the weight matrix using a scaled rotation matrix. This results in a block orthogonal matrix arrangement where, within each $2\times 2$ block, the behavior of the paired neurons exhibits spontaneous oscillations. The exploding-vanishing gradient issue can also be related to the presence of bifurcations in the parameter space of the RNN, which represent qualitative shifts in network dynamics due to parameter variations. As shown by Eisenmann et al. [224], bifurcations are associated with abrupt changes in stability regions and the topological structure of the state space. The authors of [224] propose a heuristic algorithm (SCYFI, which stands for Searcher for Cycles and Fixed points) for identifying fixed points and cycles in ReLU-based RNNs and determining their existence and stability regions in the parameter space, along with eventual bifurcations. Once that fixed points, cycles and bifurcations have been identified, Generalized Teacher Forcing (GTF) (a method aimed at redirecting diverging trajectories towards their intended targets) tends to circumvent bifurcations during training. Finally, we mention the different route followed by Echo State Networks (ESNs) and, more generally, by instances of Reservoir Computing (RC), where a large (usually non-learned) sparse connectivity layer is followed by a trainable readout function. ESN offer a compelling alternative to traditional RNNs by separating the roles of dynamic feature extraction and learning. In an ESN, the core of the network, called the reservoir, consists of a large, fixed, and sparsely connected recurrent layer, whose weights are typically randomly initialized and left untrained. This reservoir serves as a dynamic projection space that transforms input signals into a rich set of nonlinear features over time. The key innovation lies in echo state property, which ensures that the effect of any initial state fades away, making the network response dependent primarily on the recent input history. This property allows for stable dynamics without the need to train the internal connections, significantly simplifying the learning process compared to standard RNNs. Learning in ESNs is confined to the readout layer, a typically linear function that maps the high-dimensional reservoir states to the desired output. Because only this readout is trained, usually via efficient linear regression techniques, ESNs are computationally efficient and less prone to issues like vanishing or exploding gradients. The reader can find more information in [225, 226, 227].
Koopman operator view.   Koopman theory provides a powerful mathematical framework for analyzing nonlinear dynamical systems by embedding them into a linear, though typically infinite-dimensional, function space. The core idea is to focus not on the system’s state evolution itself but on how observable functions of the state evolve over time under the action of the Koopman operator, which is linear even if the underlying system is nonlinear [228]. This approach has found recent traction in machine learning, particularly for modeling time-series data, where traditional RNNs struggle with issues like vanishing gradients and lack of interpretability [56, 23]. In this context, Azencot et al. [55] propose a Consistent Koopman Autoencoder, a model that draws a direct link between Koopman theory and deep learning by embedding high-dimensional temporal data into a latent space where dynamics evolve linearly via learned Koopman operators. Unlike standard RNNs, which rely heavily on training to learn latent state transitions, this method uses physics-informed constraints, such as forward-backward consistency, to promote stability and interpretability. Their approach retains computational tractability while outperforming traditional and Koopman-based baselines, especially under noise and over longer horizons. Thus, Koopman theory not only complements RNNs by offering a principled alternative but also enriches neural forecasting architectures with physically meaningful structure and improved generalization capabilities.